OpenStudy (anonymous):

In which direction does the hyperbola given by the equation below open? (y-6)^2 over 8^2 minus (x+5)^2 over 2^2 =1

OpenStudy (mathmale):

If at all possible, type equations like this into Equation Editor. The difference in clarity is awesome.\[ (y-6)^2 \over 8^2 minus (x+5)^2 \over 2^2 =1~becomes\]

OpenStudy (mathmale):

\[\frac{ (y-6)^2 }{ 8^2 }-\frac{ (x+5)^2 }{ 2^2 }=1 \]which variable appears first in this equation, x or y? If x appears first, your hyperbola is a horizontal one, and it opens to the left and right. If y appears first, your hyp. is a vert. one, and it opens up and down.

OpenStudy (mathmale):

You'll need to find a way to remember this info for application to future problems.